OpenStudy (anonymous):

If 0.003 grams of sodium are dissolved in 100 milliliters of sweat what percentage solution is this ?

OpenStudy (kirstin012596):

There are 3 ways to find present solution:\[\frac{ weight of solute }{ volume of solution}\] \[\frac{ Volume of solute }{ volume of solution }\] \[\frac{ Weight of solute }{ Weight of solution }\] then multiply by 100 to get the percent.

OpenStudy (anonymous):

Thank you

OpenStudy (kirstin012596):

No problem